OpenStudy (anonymous):

i want to write this trig expression using power reducing identities : 6cos²x...what do I do with the 6?

OpenStudy (anonymous):

i can use the identity \[\frac{ 1+\cos2x }{ 2 }\]

OpenStudy (anonymous):

but i dont know how to deal with the 6...

OpenStudy (anonymous):

\[\cos2x = 2\cos^2x - 1\] Multiply both sides by 3 \[3\cos2x = 6 \cos^2 x - 3\]

OpenStudy (anonymous):

\[6\cos^2x = 3 \cos2x + 3\]

OpenStudy (anonymous):

ah okay!!

OpenStudy (anonymous):

@lsugano Just plug whatever you wrote above to replace cos²x, then simplify

OpenStudy (anonymous):

Wait then what if the exponent changed to like for example...4?

OpenStudy (anonymous):

like cos⁴x??

OpenStudy (anonymous):

Don't jump ahead when you don't even have the base yet!

OpenStudy (anonymous):

oh..okay

OpenStudy (anonymous):

Are you fine with this question?

OpenStudy (anonymous):

yes! i got that! i actually ran into the exact same question cos⁴x

OpenStudy (anonymous):

= (cos²x)² = ....

OpenStudy (anonymous):

cosx^4

OpenStudy (anonymous):

oh so its \[\frac{ (1+\cos2a)^2 }{ 2^2 }\]

OpenStudy (anonymous):

Yup, now expand the square :)

OpenStudy (anonymous):

\[\frac{ 1+2\cos2a+\cos^2a }{ 4 }\]

OpenStudy (anonymous):

The third term is cos ² (2a)

OpenStudy (anonymous):

ohh yess silly mistake ><

OpenStudy (anonymous):

i should bring the 4 up?

OpenStudy (anonymous):

how do i simplify than this?

OpenStudy (anonymous):

Again double the angle to reduce power!

OpenStudy (anonymous):

double?? why ?